OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.5 centimeters, what is its surface area?

OpenStudy (raden):

use the formula : surface area = 4 pi r^2 with r is radius

OpenStudy (anonymous):

whats the answer?

OpenStudy (raden):

what is 4 * 3.14 * 6.5^2 = ... ?

OpenStudy (anonymous):

530.66

OpenStudy (anonymous):

not an option

OpenStudy (anonymous):

628.57 square feet 427.43 square feet 1,335.71 square feet 854.86 square feet

OpenStudy (raden):

that's right u have to convert units cm^2 to square ft :)

OpenStudy (anonymous):

how!?

OpenStudy (raden):

that's the problem.. i dont know too try googling it :)

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

aha ok radius is half the diameter --> 1/2*6.5 = 3.25 SA = 4*pi*(3.25)^2 pi is 3.14 SA = 4(3.14)(3.25)^2 = 132.7 For surface area and volume you need to use and be familar with the formulas, thats all i was saying

OpenStudy (anonymous):

thanks can u help me some more? @nader1

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5158f430e4b0507ceba2393a @nader1